OpenStudy (anonymous):

You must answer correctly, and non-confusingly or lamas will eat your face and unicorns will give you a bad day and barf rainbows on you so you smell like leprechaun beer.

OpenStudy (anonymous):

@_@ lamas wow

OpenStudy (anonymous):

yup. I'll send Carl the human eating psychopathic lama after you.

OpenStudy (anonymous):

carl does not scare me his other half does XD

OpenStudy (anonymous):

I'll also send blue and pink unicorns over because horrible things happen when they're around

OpenStudy (anonymous):

a). 4^2 + 9^2 = c 16 + 81 = 97 \[\sqrt{97}\] i think you can solve it b). 9.8*140= is what

Parth (parthkohli):

Use pythagorean theorem: \(\Large \color{purple}{\rightarrow 9^2 + 4^2 = x^2 }\) \(\Large \color{purple}{\rightarrow 81 + 16 = x^2 }\) \(\Large \color{purple}{\rightarrow 97 = x^2 }\) New street is \(\sqrt 97\) meters.

OpenStudy (anonymous):

i went to connections too. ◕ ‿ ◕